Consider the linear transformations :

f:R^2→R^2 g:R^2→R^2

f (x,y)→(2x-y,3x+y) g(x,y)→(4x-2y,2x+y)


a) Determine the matrix of linear transformation f relative to the basis

{e1=(1,2),e2=(-1,1)}

b)dertimine (fog)(xy)

c) find g^-1


1
Expert's answer
2021-09-20T16:52:19-0400

"f(x,y)=(2x-y,\ 3x+y),\ \ \ g(x,y)=(4x-2y, \ 2x+y)"


a)Let "A" be the matrix of "f" with respect to the standard basis,

"B" be the matrix of "f" with respect to the basis "e_1=(1,2), \ e_2=(-1,1)" , and "P" be the transition matrix.

"A=\begin{pmatrix}
 2 & -1 \\
 3 & 1
\end{pmatrix}"

"P=\begin{pmatrix}
 1 & -1 \\
 2 & 1
\end{pmatrix}" , "P^{-1}=\frac{1}{3} \begin{pmatrix}
 1 & 1 \\
 -2 & 1
\end{pmatrix}=\begin{pmatrix}
 1/3 & 1/3 \\
 -2/3 & 1/3
\end{pmatrix}"

"B=P^{-1}AP =\begin{pmatrix}
 1/3 & 1/3 \\
 -2/3 & 1/3
\end{pmatrix} \begin{pmatrix}
 2 & -1 \\
 3 & 1
\end{pmatrix} \begin{pmatrix}
 1 & -1 \\
 2 & 1
\end{pmatrix}
=\begin{pmatrix}
 1/3 & 1/3 \\
 -2/3 & 1/3
\end{pmatrix} \begin{pmatrix}
 0 & -3 \\
 5 & -2
\end{pmatrix}=\begin{pmatrix}
 5/3 & -5/3 \\
 5/3 & 4/3
\end{pmatrix}"


b) "(f\circ g)(x,y) = f(4x-2y,\ 2x+y)=\big(2(4x-2y)-(2x+y),\ 3(4x-2y)+(2x+y)\big)= (6x-5y, 14x-5y)"


c) "g(x,y)=(4x-2y, \ 2x+y)=(z,w)"

Let "G" be the matrix of "f" : "G=\begin{pmatrix}
 4 & -2 \\
 2 & 1
\end{pmatrix}"

Then "G^{-1}=\frac {1}{8} \begin{pmatrix}
 1 & 2 \\
 -2 & 4
\end{pmatrix}=\begin{pmatrix}
 1/8 & 1/4 \\
 -1/4 & 1/2
\end{pmatrix}"


"g^{-1}(z,w)=(\frac{z}{8}+\frac{w}{4}, \ -\frac{z}{4}+\frac{w}{2})"


Answer:

a) "B= \begin{pmatrix}
 5/3 & -5/3 \\
 5/3 & 4/3
\end{pmatrix}"

b) "(f\circ g)(x,y) = (6x-5y, 14x-5y)"

c) "g^{-1}(z,w)=(\frac{z}{8}+\frac{w}{4}, \ -\frac{z}{4}+\frac{w}{2})"